# Dynamic Indexing Through Learned Indices with Worst-case Guarantees

### v1

@@ A.1 Finding a separating line
Algorithms 1 and 2 have a recursive depth of $O(\log n)$. Thus, if each function call takes constant time then these algorithms take $O(\log n)$ time. For ease of exposition, we showed these algorithms using three geometric predicates:
@@ B.2 Answering queries using our data structure
Given this data structure, we first show how to support member and queries (Algorithm 3). Then, we show how to answer predecessor and rank queries 5. Finally, we show how to answer range queries.
@@ B.3 Updating our data structure
For any $\varepsilon$, there exists a data structure to dynamically maintain a vertical $\varepsilon$-cover $F$ of a dynamic set of distinct integers in $O(\varepsilon+\log^{2}n+\varepsilon)$ time. We guarantee that there exists no vertical $\varepsilon$-cover $F^{\prime}$ with $|F|>\frac{3}{2}|F^{\prime}|$. The data structure supports indexing queries in $O(\varepsilon+\log|F|)$ time and range queries in additional $O(k)$ time where $k$ is the output size.
@@ Appendix C Remarks on computing a PGM-index or $\varepsilon$-cover
The degree of suboptimality depends on how strictly the problem is formulated. In its strictest sense, the output needs to be a polyline. Whenever the algorithm by O’Rourke outputs a segment $(p,q)$, the next segment needs to start with $q$. In this setting, the algorithm may output $O(n)$ lines where the optimum contains two. Indeed, consider as input the interval $[-10\varepsilon,-8\varepsilon]$ at height 00. This is followed by iterating over $i$ and adding the intervals $[-2\varepsilon,0]$ and $[0,2\varepsilon]$ at height $2i$ and $12i+$ 1, respectively. If we restrict the streaming algorithm to, after adding the segment $(p,q)$, always start at $q$ then this algorithm may output $O(n)$ lines. Indeed, the first three segments cannot be intersected by one line . If the first segment goes from $p=(-10\varepsilon,0)$ to $q=(2\varepsilon,1)$, then the next two segments cannot be intersected by a segment that starts at $q$. Repeating this process results in an output with $\frac{n}{2}+1$ segments. The optimal static solution starts with a segment from $p=(-10\varepsilon,0)$ to $q=(0,1)$ and then concludes with a segment from $q$ to $(0,2n)$. We do note that this worst-case scenario cannot occur in [13] since the segments there are $xy$-monotone.
We note that for [22] and especially [13], no such strictness is required. All that is required for [13] is a horizontal $\varepsilon$-cover (the direct horizontal analogue to our vertical $\varepsilon$-cover from Section 6). The algorithm by O’Rourke can instead also output a horizontal $\varepsilon$-cover through the same mechanism: maintain the current input segments $G$. For each insertion into $GG$, test whether there exists a segments intersecting all segments in $G$. If there does not exist such a segment, output the previous segment and delete all segments from $G$ (apart from the segment that was just inserted). This algorithm is also not optimal in a static setting, as a similar adversarial input can be constructed. However, it can easily be shown that this algorithm is a $\frac{3}{2}$ approximation.

### v2

@@ A.1 Finding a separating line
Algorithms 2 and 3 have a recursive depth of $O(\log n)$. Thus, if each function call takes constant time then these algorithms take $O(\log n)$ time. For ease of exposition, we showed these algorithms using three geometric predicates:
@@ B.2 Answering queries using our data structure
Given this data structure, we first show how to support member and queries (Algorithm 4). Then, we show how to answer predecessor and rank queries 6. Finally, we show how to answer range queries.
@@ B.3 Updating our data structure
For any $\varepsilon$, there exists a data structure to dynamically maintain a vertical $\varepsilon$-cover $F$ of a dynamic set of distinct integers in $O(\varepsilon+\log^{2}n)$ time. We guarantee that there exists no vertical $\varepsilon$-cover $F^{\prime}$ with $|F|>\frac{3}{2}|F^{\prime}|$. The data structure supports indexing queries in $O(\varepsilon+\log|F|)$ expected time and range queries in additional $O(k)$ time where $k$ is the output size.
@@ Appendix C Remarks on computing a PGM-index or $\varepsilon$-cover
The algorithm by O’Rourke can maintain a horizontal $\varepsilon$-cover in the following way: for each insertion into $G$, test whether there exists a segments intersecting all segments in $G$. If there does not exist such a segment, output the previous segment and delete all segments from $G$ (apart from the segment that was just inserted). This algorithm is also not optimal in a static setting, as a similar adversarial input can be constructed. However, it can easily be shown that this algorithm is a $\frac{3}{2}$ approximation:
@@
Appendix D Additional Experiments
We here describe our full suite of experiments, including additional real world datasets and variations of the experiments shown in Section 7. We’ve used the following datasets, each consisting of unique 8 byte integers in randomly shuffled order:
  • LINES is a synthetic data set of 5M integers that, in rank space, produces 5 lines of exponentially increasing slope. This set models the ideal scenario for a PGM index.
  • DRIFTER is a real world data set, containing roughly 1.7M steps of accumulated distance travelled by ocean drifters tracked through GPS. Used previously as a subtrajectory clustering benchmark [8].
  • BERLIN, CHICAGO, datasets each containing roughly $150.000$ GPS samples from car traffic in these cities. Used often as benchmarks for road map construction [1]. We use the accumulated distance travelled as input.
  • TIMESERIES, from [27], contains a concatenation of $15.000$ steps of time series from various real world domains. Originally used as part of the M4 forecasting competition.
D.1 Learned index complexity
As a measure of the complexity of the learned indices, we count the number of lines stored in the structures during construction. In Figure 14 the measurements across all our data can be seen. As the data arrives shuffled, even if the total data set has some underlying structure, it might not come to light until more data arrives. This is easily seen on the synthetic Lines data, where the structure starts decreasing in complexity roughly halfway for the dynamic PGM. Since the logarithmic PGM has the data split across buckets, the complexity grows until close to the end where there is a sudden drop.
Although geometric in nature, obtained as accumulated distances travelled, the Drifter, Berlin, and Chicago data sets do not show trends that indicate favorable conditions for the examined learned indices.
Figure 14: The number of lines stored in the learned indices, insertion only.
D.2 Maintaining an indexing data structure
Here we show the full results from varying query ratio and prior deletion ratio on all of the datasets described above.
The learned index.
When measuring the running times of maintaining the learned index, the key observation is that the smaller structured datasets Drifter, Berlin, and Chicago have much closer performance than the larger Longitude data. As is the case also on the unstructured data, the performance penalties incurred by the poor memory access patterns of the convex hull data structures becomes increasingly prohibitive once data surpasses a certain size.
Figure 15: Update times for maintaining the learned index dynamically.
The indexing data structure.
For the larger datasets, Longitude, Unif, and Book, we see some interesting trends. Longitude is the only structured large set, where we see an increase of performance from the dynamic PGM as query ratio increases, whereas it decreases for the logarithmic PGM. In the unstructured cases, both structures occur more time spent for larger query ratios. For the smaller datasets, performance is almost identical. The overhead associated with maintaining the learned index becomes much smaller, since it can now be manipulated with less bad effects on cache. Still, memory access is the deciding factor, and for many cases the logarithmic PGM comes out on top.
As we increase the number of previous deletions, the gap in performance between indexing structures lessens. This somewhat ties into the previous notes on memory, but we also see that for the larger query ratios performance evens out, with the dynamic PGM only a sliver from the logarithmic PGM on structured data.
The adversarial scenario is taking the above to its extreme, and the dynamic then outperforms the logarithmic on all data.
Figure 16: Time per operation in dynamic scenario with varying query ratio and prior deletions on Longitude data
Figure 17: Time per operation in dynamic scenario with varying query ratio and prior deletions on Unif data
Figure 18: Time per operation in dynamic scenario with varying query ratio and prior deletions on Drifter data
Figure 19: Time per operation in dynamic scenario with varying query ratio and prior deletions on Book data
Figure 20: Time per operation in dynamic scenario with varying query ratio and prior deletions on Berlin data
Figure 21: Time per operation in dynamic scenario with varying query ratio and prior deletions on Chicago data
Figure 22: Time per operation in dynamic scenario with varying query ratio and prior deletions on Timeseries data